OpenStudy (anonymous):

A room in the shape of a triangle has sides of length 7 yd, 8 yd, and 11 yd. If carpeting costs $17.50 a square yard and padding costs $3.25 a square yard, how much to the nearest dollar will it cost to carpet the room, assuming that there is no waste?

OpenStudy (anonymous):

i think i figured this out at $579.77 0r $580 :)

OpenStudy (anonymous):



OpenStudy (anonymous):

i got $581.00. soo idk.. lol

OpenStudy (anonymous):

Using heron's forumula for calculating the height s=(a+b+c)/2 h=2/c *sqrt(s(s-a)(s-b)(s-c)) Plugging in a=8, b=7, c=11, I get h=5.08 Then the area is 5.08 * 11 = 55.86 Multiply that by the total cost of materials: cost= 55.86*(3.25+17.50) I get $1159.03

OpenStudy (anonymous):

oops, and divide by 2 that's 579.52

OpenStudy (anonymous):

\[A=1/2\times b \times h\] \[A=(.5) \times (7) \times (8) \] = 28SQ. Feet $17.50 x 28= $490 $3.25 x 28+ $91 $490+91=$581.00

OpenStudy (anonymous):

emistre: 7 and 8 are not the base and height of the triangle. It's not a right triangle.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The answer was 580 as we were to round up. Thank you for the assistance!